# Assessing Inf luential Observations in Pain Prediction using fMRI Data

### v1

@@ 3.2 Multiple Inf luential Observations Detection
To detect multiple influential observations, we first review the multiple influential point (MIP) detection algorithm by Zhao et al. (2019) in Section 3.2.1, which is an extension of the HIM procedure discussed in Section 3.1.1. This is followed by the proposal of our clustering-based multiple influential point (ClusMIP) detection method in Section 3.2.2.
@@ 3.2.1 Multiple Inf luential Point (MIP) Detection Procedure
The leave-one-out basis of the HIM approach implies that it may not be effective in capturing multiple influential points. To remedy this issue, Zhao et al. (2019) propose the MIP method via integrating the so-called random-group-deletion (RGD) scheme into the HIM framework. Based on this, two HIM-based statistics are subsequently formulated to overcome both masking and swamping effects in detecting multiple influential observations.
To have a proper understanding of this procedure, we first explain the philosophy underlying the RGD algorithm, which is provided in Section 5 of the Supplementary Materials. The RDG scheme centers on seeking and subsequently attaching a clean subset of data to every single point, of which refinement is conducted based on the combined dataset. To be more precise, for each observation $i$, $m$ samples indexed by $A_{1},...,A_{m}$ are drawn uniformly and randomly from the remaining portion of the data. When $m$ is sufficiently large, there is a high probability that the subset $\{\text{Z}_{r}:r\in A_{j}\}$ is clean for some $j\in\{1,...,m\}$. Thus, if $i^{\text{th}}$ observation is contaminated, it would be the sole influential point in the merged subset $\{\text{Z}_{r}:r\in A_{j}\cup\{i\}\}$. Thus, further assessment can be carried out on this combined dataset.
Toward that end, for each $i$, two HIM-based statistics $\text{T}_{\text{min},i,m}$ and $\text{T}_{\text{max},i,m}$ are then constructed, which are respectively defined by $\text{T}_{\text{min},i,m}=\arg\!\min_{1\leq j\leq m}n_{\text{sub}}^{2}D_{j,i}$ and $\text{T}_{\text{max},i,m}=\mathop{\mathrm{argmax}}_{1\leq j\leq m}n_{\text{sub}}^{2}D_{j,i}$, where $D_{j,i}=\sum_{k=1}^{p}\big{(}\widehat{\rho}_{k}^{(A_{j}\cup\{i\})}-\widehat{\rho}_{k}^{(A_{j}}\big{)}^{2}/p$. Here, $\widehat{\rho}_{k}^{(A_{j}\cup\{i\})}$ and $\widehat{\rho}_{k}^{(A_{j})}$ are the sample marginal correlations between the response Y and $k^{\text{th}}$ predictor $x_{k}$ based on $\{\text{Z}_{r}:r\in\{A_{j}\cup\{i\}\}\}$ and $\{\text{Z}_{r}:r\in A_{j}\}$. The authors then establish that they are both asymptotically $\chi^{2}(1)$ and indeed complement each other: $\text{T}_{\text{min},i,m}$ tends to select conspicuously influential points. Thus, it is effective in overcoming the swamping effect. However, it may be conservative so that true influential points may be neglected. In contrast, $\text{T}_{\text{max},i,m}$ is powerful to counter the masking effect. Yet, it may be inordinately aggressive so that clean ones may be falsely chosen. Therefore, combining the strengths of both $\text{T}_{\text{min},i,m}$ and $\text{T}_{\text{max},i,m}$ leads to the MIP detection procedure provided in Section 6 of the Supplementary Materials. In practice, the MIP algorithm is implemented in the $\tt{R}$ package $\tt{MIP}$.
@@ 3.2.2 Clustering-based MIP (ClusMIP) Detection Procedure
Motivated by Zhao et al. (2015, 2019), one apparent solution to detect multiple influential observations specifically on variable selection is to integrate the aforementioned RGD scheme into the GDF framework (Section 3.1.3), leading to the proposal of the RGD-based GDF (RGDF) procedure provided in Section 7 of the Supplementary Material.
@@
To alleviate the exorbitant computational costs, instead of the foregoing RGD strategy, we incorporate an appropriate high-dimensional clustering procedure for initially partitioning the complete dataset into approximately clean ($\widehat{\text{S}}_{\text{clean}}$) and influential ($\widehat{\text{S}}_{\text{infl}}$) proportions. Then, each observation $k$ in $\widehat{\text{S}}_{\text{infl}}$ is further assessed based on the sole subset of the data indexed by $\{k\}\cup\widehat{\text{S}}_{\text{clean}}$ via Algorithm 1 which is proposed for detecting a single influential observation through the GDF measure (5). In addition, to simultaneously test the $\rvert\widehat{\text{S}}_{\text{infl}}\lvert$ hypotheses that the $k^{\text{th}}$ observation is not influential, $k\in\widehat{\text{S}}_{\text{infl}}$, the procedure of Benjamini & Hochberg (1995) is adopted to address the multiple testing issue. This leads to our ClusMIP algorithm summarized in Algorithm 2.
Algorithm 2 Clustering-based Multiple Influential Point Detection (ClusMIP)
@@
2:Given a nominal level $\alpha\in(0,1)$, the final estimated influential observations set is $\widehat{\text{I}}_{\text{infl}}=\{j\in\widehat{\text{S}}_{\text{infl}}:\lvert(\tau_{j}-\tilde{\tau}_{j})/\tilde{\sigma}_{\tau_{j}}\rvert>\gamma_{\alpha}\}$, where for each $j\in\widehat{\text{S}}_{\text{infl}}$, $\tau_{j}$ is computed from subset indexed by $\{j\}\cup\widehat{\text{S}}_{\text{clean}}$, and $\tilde{\tau}_{j}$ and $\tilde{\sigma}_{\tau_{j}}$ are the sample mean and variance of $\{\tau_{k}:k\in\{j\}\cup\widehat{\text{S}}_{\text{clean}}\}$. To simultaneously test hypotheses that $k^{\text{th}}$ point is not influential, $k\in\widehat{\text{S}}_{\text{infl}}$, procedure of Benjamini & Hochberg (1995) is adopted.
@@
Toward that end, one possible approach is $K$-means clustering (KC), which minimizes the within-cluster Euclidean distance. Pollard (1981) provides conditions for the strong consistency of the KC, showing that with the correct number of clusters, the cluster centers converge to their true counterparts as $n\rightarrow\infty$ under certain conditions. Yet, this approach is limited on three facets. First, its widely adopted approximating algorithm, known as Lloyd’s algorithm, may not be globally optimal due to its heavy sensitivity to the starting points. Second, the consistency of the KC may not hold under diverging dimension $p\rightarrow\infty$. Third, under such high-dimensional settings, the notion of distance in terms of the $\ell^{p}$ norm may not be meaningful (Beyer et al., 1999). Indeed, it is justified [ibid., Theorem 1] that as $p\rightarrow\infty$, the ratio of the distances in $\ell^{p}$ norm of the nearest and furthest points to a reference point approaches 1, implying that points tend to be uniformly separated in the high-dimensional $\ell^{p}$ space. In fact, this assertion aligns with the geometric characterization of high-dimensional data (Hall et al., 2005), showing that as $p$ increases while $n$ remains fixed, the data tend to cluster at the vertices of a deterministic simplex.
@@
To address the first limitation, an improved choice of the initial values for the KC algorithm is proposed by Arthur & Vassilvitskii (2007), leading to the so-called $K$-means++ algorithm (KC++). In addition, to resolve the challenges brought about by the high dimensionality, two categories of remedial proposals are available in the literature, which are known as the projection-based and regularized clustering. The projection-based scheme is a hybrid technique by first projecting the high-dimensional data onto a low-dimensional space, followed by partition via the KC. Two proposals under this category are the $t$-distributed stochastic neighbor embedding ($t$-SNE) (Hinton & Roweis, 2002) and spectral clustering (Ng et al., 2001), where the projected low-dimensional space for the later work is spanned by eigenvectors of the normalized graph Laplacians. In comparison, the regularized clustering augments a penalty term to the appropriate objective function to obtain estimates of penalized cluster centers. One such approach is called regularized KC (Sun et al., 2012), where a group LASSO penalty term is adopted and strong consistency of cluster centers is established [ibid., Theorem 1] for $p\rightarrow\infty$. These five clustering procedures, including the KC, KC++, $t$-SNE, spectral clustering and regularized KC are empirically compared on all the simulation configurations considered in Section 4. The corresponding graphical illustrations are reported in Section 9 of the Supplementary Materials, where all of them exhibit analogously satisfactory detection of clusters. Indeed, in both finite and diverging dimensions, with a consistent clustering technique that is further refined via the CLT of the GDF measure (5) established in Theorem 2, the FPR of the foregoing ClusMIP algorithm is controlled at a given the FDR which is stated below, where the proof is given in Section 8 of the Supplementary Materials:
@@
Suppose that the proportion of the true clean observations is more than 1/2. Then, given the FDR controlled at $\alpha_{0}\in(0,1)$, we have
@@
    $\mathbb{E}[\text{FPR}(\text{ClusMIP})]\leq\alpha_{0}$ at any finite dimensions $n$ and $p$, and
@@
    $\mathbb{E}[\text{FPR}(\text{ClusMIP})]\rightarrow 0$ as $\min{(n,p)}\rightarrow\infty$.
where the FPR(ClusMIP) refers to the FPR of the ClusMIP procedure given in Algorithm 2.
@@ 4 Simulation Study
We now conduct simulation study to assess and compare the performance of the various detection techniques discussed in Section 3 for high-dimensional data. The details of data generation are provided in Section 4.1, followed by evaluation criteria and implementation details respectively in Sections 4.2 and 4.3. Analysis of results is given in Section 4.4.
@@ 4.1 Data Generation
Let $(n,p)=(100,1000)$. X follows a multivariate normal distribution, generated as $\textbf{X}\sim\mathcal{N}(\boldsymbol{0},\Sigma(\rho))$, where $\Sigma(\rho)$ is the Toeplitz-type variance-covariance matrix indexed by the correlation coefficient $\rho$, and its $(i,j)$-entry is given by $\Sigma_{i,j}(\rho)=\rho^{|i-j|}$, $i=1,...,p$ and $j=1,...,p$. Here, we consider $\rho=0,0.5$ and $0.8$ so that the corresponding design matrices are denoted by $\textbf{X}_{\text{Tp(0)}}$, $\textbf{X}_{\text{Tp(0.5)}}$ and $\textbf{X}_{\text{Tp(0.8)}}$, respectively.
@@
We consider the data generating mechanism via the linear regression model (1). Specifically, we simulate influential and clean observations according to the following two-step procedure: first, generate $n=100$ observations based on (1). Second, generate $n_{\text{infl}}$ points, denoted by $\{\textbf{Z}_{i,\text{infl}}=(\text{Y}_{i,\text{infl}},\textbf{X}_{i,\text{infl}}^{\top})^{\top}\in\mathbb{R}^{1001}:i=1,...,n_{\text{infl}}\}$, according to three perturbation models presented below, where $n_{\text{infl}}$ is the number of influential points. The first $n_{\text{infl}}$ points in the first step are then replaced by $\textbf{Z}_{i,\text{infl}},i=1,..,n_{\text{infl}}$, yielding the heterogeneously contaminated dataset. Here, the three perturbation models respectively introduce contamination to the response vector only, design matrix only and both of them simultaneously. Furthermore, the proportion of contamination $\zeta\coloneqq(n_{\text{infl}}/n)\times 100\%$ is considered to be $5\%,10\%,15\%$ and $20\%$. From this two-step scheme, we next provide data generation specifics for (1), which are also summarized in Section 10.1 of the Supplementary Materials.
Linear Regression Model
The true regression coefficient in (1) is $\boldsymbol{\beta}=\{2,1,1,0,...,0\}^{\top}\in\mathbb{R}^{p}$ and $\sigma^{2}=1$. To generate influential points, we define $i_{\text{linear},\text{max}}=\mathop{\mathrm{argmax}}_{1\leq i\leq n}\text{Y}_{i}$ to facilitate the distinction of them from the clean portion based on the chosen clustering procedure. Then, for $i=1,...,n_{\text{infl}}$, we adapt and modify the settings in Zhao et al. (2013) and propose:
@@
    Perturbation Model I (Contamination on the Response): $\textbf{X}_{i,\text{infl}}=\textbf{X}_{i}$ and $\text{Y}_{i,\text{infl}}=\text{Y}_{i_{\text{linear},\text{max}}}\\ +\epsilon_{i}+\kappa_{\text{linear},\textbf{Y}}$, where $\epsilon_{i}\sim N(0,1)$ and $\kappa_{\text{linear},\textbf{Y}}\in\{5,10\}$.
@@
    Perturbation Model II (Contamination on the Predictors): $\text{X}_{ij,\text{infl}}=\text{X}_{ij}+\kappa_{\text{linear},\textbf{X}}\cdot\mathbbm{1}\big{\{}j\in\{1,...,10\}\big{\}}$, $\kappa_{\text{linear},\textbf{X}}\in\{5,10\}$, $\textbf{X}_{i,\text{infl}}=(\text{X}_{i1,\text{infl}},...,\text{X}_{ip,\text{infl}})^{\top}$, and $\text{Y}_{i,\text{infl}}=\text{Y}_{i}$.
@@
    Perturbation Model III (Contamination on the Response and Predictors): $\text{X}_{ij,\text{infl}}=\\ \text{X}_{i_{\text{linear},\text{max}}j}+0.5\ \kappa_{\text{linear},\textbf{X}}\cdot\mathbbm{1}\big{\{}j\in\{1,...,10\}\big{\}}$, $\textbf{X}_{i,\text{infl}}=(\text{X}_{i1,\text{infl}},...,\text{X}_{ip,\text{infl}})^{\top}$, and $\text{Y}_{i,\text{infl}}=(\textbf{X}_{i,\text{infl}})^{\top}\boldsymbol{\beta}+\epsilon_{i}+\kappa_{\text{linear},\textbf{Y}}$, where $\epsilon_{i}\sim N(0,1)$ and $(\kappa_{\text{linear},\textbf{X}},\kappa_{\text{linear},\textbf{Y}})\in\{(5,5),(10,10)\}$.

### v2

@@ 4.2 Multiple Inf luential Observations Detection
To detect multiple influential observations, we first review the multiple influential point (MIP) detection algorithm by [39] in Section 4.2.1, which is an extension of the HIM procedure discussed in Section 4.1.1. This is followed by the proposal of our clustering-based multiple influential point (ClusMIP) detection method in Section 4.2.2.
@@ 4.2.1 Multiple Inf luential Point (MIP) Detection Procedure
The leave-one-out basis of the HIM approach implies that it may not be effective in capturing multiple influential points. To remedy this issue, [39] propose the MIP method via integrating the so-called random-group-deletion (RGD) scheme into the HIM framework. Based on this, two HIM-based statistics are subsequently formulated to overcome both masking and swamping effects in detecting multiple influential observations.
To have a proper understanding of this procedure, we first explain the philosophy underlying the RGD algorithm, which is provided in Section 6 of the Supplementary Materials Part I. The RDG scheme centers on seeking and subsequently attaching a clean subset of data to every single point, of which refinement is conducted based on the combined dataset. To be more precise, for each observation $i$, $m$ samples indexed by $A_{1},...,A_{m}$ are drawn uniformly and randomly from the remaining portion of the data. When $m$ is sufficiently large, there is a high probability that the subset $\{\text{Z}_{r}:r\in A_{j}\}$ is clean for some $j\in\{1,...,m\}$. Thus, if $i^{\text{th}}$ observation is contaminated, it would be the sole influential point in the merged subset $\{\text{Z}_{r}:r\in A_{j}\cup\{i\}\}$. Thus, further assessment can be carried out on this combined dataset.
Toward that end, for each $i$, two HIM-based statistics $\text{T}_{\text{min},i,m}$ and $\text{T}_{\text{max},i,m}$ are then constructed, which are respectively defined by $\text{T}_{\text{min},i,m}=\arg\!\min_{1\leq j\leq m}n_{\text{sub}}^{2}D_{j,i}$ and $\text{T}_{\text{max},i,m}=\mathop{\mathrm{argmax}}_{1\leq j\leq m}n_{\text{sub}}^{2}D_{j,i}$, where $D_{j,i}=\sum_{k=1}^{p}\big{(}\widehat{\rho}_{k}^{(A_{j}\cup\{i\})}-\widehat{\rho}_{k}^{(A_{j}}\big{)}^{2}/p$. Here, $\widehat{\rho}_{k}^{(A_{j}\cup\{i\})}$ and $\widehat{\rho}_{k}^{(A_{j})}$ are the sample marginal correlations between the response Y and $k^{\text{th}}$ predictor $x_{k}$ based on $\{\text{Z}_{r}:r\in\{A_{j}\cup\{i\}\}\}$ and $\{\text{Z}_{r}:r\in A_{j}\}$. The authors then establish that they are both asymptotically $\chi^{2}(1)$ and indeed complement each other: $\text{T}_{\text{min},i,m}$ tends to select conspicuously influential points. Thus, it is effective in overcoming the swamping effect. However, it may be conservative so that true influential points may be neglected. In contrast, $\text{T}_{\text{max},i,m}$ is powerful to counter the masking effect. Yet, it may be inordinately aggressive so that clean ones may be falsely chosen. Therefore, combining the strengths of both $\text{T}_{\text{min},i,m}$ and $\text{T}_{\text{max},i,m}$ leads to the MIP detection procedure provided in Section 7 of the Supplementary Materials Part I. In practice, the MIP algorithm is implemented in the $\tt{R}$ package $\tt{MIP}$.
@@ 4.2.2 Clustering-based MIP (ClusMIP) Detection Procedure
Motivated by [38, 39], one apparent solution to detect multiple influential observations on variable selection is to integrate the RGD scheme into the GDF metric (3), leading to the RGD-based GDF (RGDF) procedure provided in Section 8 of the Supplementary Materials Part I.
@@
To alleviate the computational costs, instead of the RGD strategy, we incorporate an appropriate high-dimensional clustering procedure for initially partitioning the complete dataset into approximately clean ($\widehat{\text{S}}_{\text{clean}}$) and influential ($\widehat{\text{S}}_{\text{infl}}$) portions. Then, each point $k$ in $\widehat{\text{S}}_{\text{infl}}$ is further assessed solely based on the subset of the data indexed by $\{k\}\cup\widehat{\text{S}}_{\text{clean}}$ via Algorithm GDF which is proposed for detecting a single influential point through the GDF measure (3). Moreover, to simultaneously test the $\rvert\widehat{\text{S}}_{\text{infl}}\lvert$ hypotheses that the $k^{\text{th}}$ point is clean, $k\in\widehat{\text{S}}_{\text{infl}}$, the procedure of [4] is adopted to address the multiple testing issue. This leads to our ClusMIP procedure summarized below in Algorithm ClusMIP.
Algorithm ClusMIP Clustering-based Multiple Influential Point Detection (ClusMIP)
@@
2:Given a nominal level $\alpha\in(0,1)$, the final estimated influential observations set is $\widehat{\text{I}}_{\text{infl}}=\{j\in\widehat{\text{S}}_{\text{infl}}:\lvert(\tau_{j}-\tilde{\tau}_{j})/\tilde{\sigma}_{\tau_{j}}\rvert>\gamma_{\alpha}\}$, where for each $j\in\widehat{\text{S}}_{\text{infl}}$, $\tau_{j}$ defined in (3) is computed from subset indexed by $\{j\}\cup\widehat{\text{S}}_{\text{clean}}$, and $\tilde{\tau}_{j}$ and $\tilde{\sigma}_{\tau_{j}}$ are the sample mean and variance of $\{\tau_{k}:k\in\{j\}\cup\widehat{\text{S}}_{\text{clean}}\}$. Here, $\gamma_{\alpha}$ is the $100\cdot(1-\alpha/2)^{\text{th}}$ percentile of a standard normal $N(0,1)$. To simultaneously test hypotheses that $k^{\text{th}}$ point is not influential, $k\in\widehat{\text{S}}_{\text{infl}}$, procedure of [4] is adopted.
@@
Toward that end, one possible approach is $K$-means clustering (KC), which minimizes the within-cluster Euclidean distance. [27] provides conditions for the strong consistency of the KC, showing that with the correct number of clusters, the cluster centers converge to their true counterparts as $n\rightarrow\infty$ under certain conditions. Yet, this approach is limited on three facets. First, its widely adopted approximating algorithm, known as Lloyd’s algorithm, may not be globally optimal due to its heavy sensitivity to the starting points. Second, the consistency of the KC may not hold under diverging dimension $p\rightarrow\infty$. Third, under such high-dimensional settings, the notion of distance in terms of the $\ell^{p}$ norm may not be meaningful [5]. Indeed, it is justified [ibid., Theorem 1] that as $p\rightarrow\infty$, the ratio of the distances in $\ell^{p}$ norm of the nearest and furthest points to a reference point approaches 1, implying that points tend to be uniformly separated in the high-dimensional $\ell^{p}$ space.
@@
To address the first limitation, an improved choice of the initial values for the KC algorithm is proposed by [1], leading to the so-called $K$-means++ algorithm (KC++). In addition, to resolve the challenges brought about by the high dimensionality, two categories of remedial proposals are available in the literature, which are known as the projection-based and regularized clustering. The projection-based scheme is a hybrid technique by first projecting the high-dimensional data onto a low-dimensional space, followed by partition via the KC. Two proposals under this category are the $t$-distributed stochastic neighbor embedding ($t$-SNE) [18] and spectral clustering [24], where the projected low-dimensional space for the later work is spanned by eigenvectors of the normalized graph Laplacians. In comparison, the regularized clustering augments a penalty term to the appropriate objective function to obtain estimates of penalized cluster centers. One such approach is called regularized KC [30], where a group LASSO penalty term is adopted and strong consistency of cluster centers is established [ibid., Theorem 1] for $p\rightarrow\infty$. These five clustering procedures, including the KC, KC++, $t$-SNE, spectral clustering and regularized KC are empirically compared on all the simulation configurations considered in Section 5. The corresponding graphical illustrations are reported in Section 10 of the Supplementary Materials Part I, where all of them exhibit analogously satisfactory detection of clusters. Indeed, in both finite and diverging dimensions, with a consistent clustering technique that is further refined via the CLT of the GDF measure (3) established in Theorem 2, the FPR of the foregoing ClusMIP algorithm is controlled at a given the FDR which is stated below in Theorem 3. The consistency conditions (B.1) to (B.2) [27], (C.1) to (C.7) [29], and the proof are given in Section 9 of the Supplementary Materials Part I:
@@
Suppose that the proportion of clean observations is more than 1/2 and FDR is controlled at $\alpha_{0}\in(0,1)$. Then,
@@
    $\mathbb{E}[\text{FPR}(\text{ClusMIP})]\leq\alpha_{0}$ at any finite dimensions $n$ and $p$.
@@
    If $p$ is fixed, then $\mathbb{E}[\text{FPR}(\text{ClusMIP})]\rightarrow 0$ as $n\rightarrow\infty$ provided that (B.1) and (B.2) hold.
  3. (c)
    If $p\rightarrow\infty$, then $\mathbb{E}[\text{FPR}(\text{ClusMIP})]\rightarrow 0$ as $n\rightarrow\infty$ provided that (C.1) to (C.7) hold.
where the FPR(ClusMIP) refers to the FPR of the ClusMIP given in Algorithm ClusMIP.
The aforementioned conditions (B.1) to (B.2) and (C.1) to (C.7) are standard assumptions in clustering. They ensure the uniqueness of the true cluster centers. Moreover, they require that the second moment of the distribution of data points is finite, and the distribution is continuous. Condition (C.5) is a technical condition requiring the existence of a spherically symmetric envelop for the underlying density whose $p^{\text{th}}$ moment exists. This assumption requires some control on the tail behavior of the underlying density. It essentially says that the tails of the density of data points should vanish polynomially, of degree $p$ fast. As a result, the higher the dimension of the space, $p$, the faster the tails should die out.
@@ 5 Simulation Study
The data generation is explained in Section 5.1, followed by evaluation metric and implementation in Sections 5.2 and 5.3. Analysis of results is in Section 5.4.
@@ 5.1 Data Generation
Let $(n,p)=(100,1000)$ and $(160,480)$, where the latter dimension setting resembles those obsereved in the motivating fMRI study [21]. For each set of dimensions, $\textbf{X}\in\mathbb{R}^{n\times p}$ follows a zero-mean multivariate normal distribution, denoted $\textbf{X}\sim\mathcal{N}(\boldsymbol{0},\Sigma)$. Two possible variance-covariance matrices $\Sigma\in\mathbb{R}^{p\times p}$ are considered: the Toeplitz and the spatially correlated structures. Specifically, their $(i,j)$-entry is given by $\rho_{1}^{|i-j|}$ and $\rho_{2}^{|i-j|}\mathbbm{1}(i,j\in\{(p-S)/2+1,...,(p-S)/2+S\})$, $i=1,...,p$ and $j=1,...,p$, respectively. The spatially correlated formulation, where only the center $S$ predictors are dependent, mimics the brain connectivity patterns observed in fMRI. In our simulation study, we let $\rho_{1}\in\{0,0.5,0.8\}$, $\rho_{2}=0.8$ and $S=100$ so that four design matrices respectively denoted by $\textbf{X}_{\text{Tp(0)}}$, $\textbf{X}_{\text{Tp(0.5)}}$, $\textbf{X}_{\text{Tp(0.8)}}$ and $\textbf{X}_{\text{Sc(0.8)}}$ are generated for each dimension.
@@
Influential and clean observations are synthesized using the following two-step procedure: first, generate $n=100$ observations based on (1). Second, generate $n_{\text{infl}}$ points, denoted by $\{\textbf{Z}_{i,\text{infl}}=(\text{Y}_{i,\text{infl}},\textbf{X}_{i,\text{infl}}^{\top})^{\top}\in\mathbb{R}^{1001}:i=1,...,n_{\text{infl}}\}$, according to three perturbation schemes presented below, where $n_{\text{infl}}$ is the number of influential points. The first $n_{\text{infl}}$ points in the first step are then replaced by $\textbf{Z}_{i,\text{infl}},i=1,..,n_{\text{infl}}$, yielding the heterogeneous, contaminated dataset. Here, the three schemes respectively introduce contamination to the response vector only, design matrix only and both of them simultaneously. Moreover, the proportion of contamination $\zeta\coloneqq(n_{\text{infl}}/n)\times 100\%$ is set to be $5\%,10\%,15\%$ and $20\%$. The data generation is given in Section 11.1 of the Supplementary Materials Part I.
Perturbation Schemes
The true regression coefficient in (1) is $\boldsymbol{\beta}=\{2,1,1,0,...,0\}^{\top}\in\mathbb{R}^{p}$ and $\sigma^{2}=1$. To generate influential points, we define $i_{\text{linear},\text{max}}=\mathop{\mathrm{argmax}}_{1\leq i\leq n}\text{Y}_{i}$ to facilitate the distinction of them from the clean portion based on the chosen clustering method. Then, for $i=1,...,n_{\text{infl}}$, we adapt and modify the settings in [37] and propose the following three schemes each with low, moderate and high contamination intensities/scales:
@@
    Scheme I (Contamination on the Response): $\textbf{X}_{i,\text{infl}}=\textbf{X}_{i}$ and $\text{Y}_{i,\text{infl}}=\text{Y}_{i_{\text{linear},\text{max}}}\\ +\epsilon_{i}+\kappa_{\text{linear},\textbf{Y}}$, where $\epsilon_{i}\sim N(0,1)$ and $\kappa_{\text{linear},\textbf{Y}}\in\{5,10,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}20}\}$.
@@
    Scheme II (Contamination on the Predictors): $\text{X}_{ij,\text{infl}}=\text{X}_{ij}+\kappa_{\text{linear},\textbf{X}}\cdot\mathbbm{1}\big{\{}j\in\{1,...,10\}\big{\}}$, $\kappa_{\text{linear},\textbf{X}}\in\{5,10,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}20}\}$, $\textbf{X}_{i,\text{infl}}=(\text{X}_{i1,\text{infl}},...,\text{X}_{ip,\text{infl}})^{\top}$, and $\text{Y}_{i,\text{infl}}=\text{Y}_{i}$.
@@
    Scheme III (Contamination on the Response and Predictors): $\text{X}_{ij,\text{infl}}=\\ \text{X}_{i_{\text{linear},\text{max}}j}+0.5\ \kappa_{\text{linear},\textbf{X}}\cdot\mathbbm{1}\big{\{}j\in\{1,...,10\}\big{\}}$, $\textbf{X}_{i,\text{infl}}=(\text{X}_{i1,\text{infl}},...,\text{X}_{ip,\text{infl}})^{\top}$, and $\text{Y}_{i,\text{infl}}=(\textbf{X}_{i,\text{infl}})^{\top}\boldsymbol{\beta}+\epsilon_{i}+\kappa_{\text{linear},\textbf{Y}}$, where $\epsilon_{i}\sim N(0,1)$ and $(\kappa_{\text{linear},\textbf{X}},\kappa_{\text{linear},\textbf{Y}})\in\{(5,5),(10,10),{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(20,20)}\}$.